# Towards Context-Aware Image Anonymization with Multi-Agent Reasoning

### 29 Mar 2026 v1

@@ Abstract
Street-level imagery contains personally identifiable information (PII), some of which is context-dependent. Existing anonymization methods either over-process images or miss subtle identifiers, while API-based solutions compromise data sovereignty. We present an agentic framework CAIAMAR (Context-Aware Image Anonymization with Multi-Agent Reasoning) for context-aware PII segmentation with diffusion-based anonymization, combining pre-defined processing for high-confidence cases with multi-agent reasoning for indirect identifiers. Three specialized agents coordinate via round-robin speaker selection in a Plan-Do-Check-Act (PDCA) cycle, enabling large vision-language models to classify PII based on spatial context (private vs. public property) rather than rigid category rules. The agents implement spatially-filtered coarse-to-fine detection where a scout-and-zoom strategy identifies candidates, open-vocabulary segmentation processes localized crops, and IoUIoU-based deduplication ($30\%$ threshold) prevents redundant processing. Modal-specific diffusion guidance with appearance decorrelation substantially reduces re-identification (Re-ID) risks. On CUHK03-NP, our method reduces person Re-ID risk by $73\%$ ($R1$: $16.9\%$ vs. $62.4\%$ baseline). For image quality preservation on CityScapes, we achieve KID: $0.001$, and FID: $9.1$, significantly outperforming existing anonymization. The agentic workflow detects non-direct PII instances across object categories, and downstream semantic segmentation is preserved. Operating entirely on-premise with open-source models, the framework generates human-interpretable audit trails supporting EU’s GDPR transparency requirements while flagging failed cases for human review.
@@ B.1 Obfuscation and Generative Methods
Conventional image anonymization approaches, such as EgoBlur [30], rely on blurring for efficiency. However, these methods insufficiently protect identity information while degrading downstream task performance. For instance, a $5.3\%$ drop in instance segmentation average precision (APAP) on CityScapes [5] has been reported when applying Gaussian blur [13]. Moreover, blurring remains vulnerable to inversion attacks: $95.9\%$ identity retrieval accuracy has been achieved on blurred CelebA-HQ [15] face images at $256\times 256$ resolution [43], demonstrating that simple obfuscation is inadequate for modern threat models. Generative methods address these limitations by synthesizing realistic replacements for sensitive regions, preserving contextual consistency while reducing Re-ID risks. Early deep learning–based methods predominantly use Generative Adversarial Networks (GANs). DeepPrivacy [14] employs a conditional GAN that generates anonymized faces conditioned on pose keypoints and background context. DeepPrivacy2 [12] extends this approach to full-body anonymization via a style-based GAN, improving realism and consistency across human appearance and pose. Recent advances leverage diffusion models, offering higher fidelity, controllability, and semantic alignment. LDFA [16] applies latent diffusion for face anonymization in street recordings, maintaining both realism and privacy. The FADM framework [47] and RAD [25] employ diffusion to anonymize persons. SVIA [21] extends diffusion-based anonymization to large-scale street-view imagery, ensuring privacy preservation in autonomous driving datasets by replacing the semantic categories of person, vehicle, traffic sign, road, and building. Reverse Personalization [17] introduces a finetuning-free face anonymization framework based on conditional diffusion inversion with null-identity embeddings. By applying negative classifier-free guidance, it successfully steers the generation process away from identity-defining features while enabling attribute-controllable anonymization for arbitrary subjects.
@@ C.2 Phase 1: Pre-defined Preprocessing
We employ YOLOv8m-seg [^1] https://github.com/ultralytics/ultralytics, Accessed October 25, 2025 for instance segmentation (confidence threshold $\tau=0.25$). The anonymization pipeline operates in three stages: (1) extracts visual descriptions via Qwen2.5-VL-32B [2], capturing pose, body build, view angle, and actions; (2) generates anonymized prompts by instructing the LVLM to randomly select clothing colors and brightness modifiers from predefined palettes (2020 colors, $10$ brightness levels) [^2] Color palette: gray, beige, navy, white, black, brown, khaki, blue, red, green, yellow, pink, teal, burgundy, olive, charcoal, maroon, tan, cream, sage. Brightness modifiers: light, dark, bright, faded, vibrant, muted, pale, deep, pastel, bold. rather than describing observed clothing, while preserving utility-relevant attributes (body build, pose, view angle, actions) and uniformed personnel context; (3) performs SDXL inpainting at $768$ px resolution with OpenPose [3] ControlNet333thibaud/controlnet-openpose-sdxl-1.0 (conditioning scale $\alpha=0.8$, strength $\rho=0.9$, guidance scale $s=9.0$, $25$ denoising steps). The positive and negative prompts (excluding the LVLM-generated description) are adapted from [47]. Critically, we disable all color matching to prevent appearance correlation and eliminate Re-ID vectors.
License Plate Processing. Small object detection challenges necessitate specialized handling. We train YOLOv8s on UC3M-LP dataset [31] at 12801280px resolution, achieving $mAP_{50-95}=0.82$. At inference, we apply reduced confidence threshold ($\tau=0.05$) prioritizing recall and Non-Maximum Suppression (NMS) with IoUIoU threshold of $0.5$ to ensure individual plate separation, aligning with conventional NMS practices [37]. Gauss. blur (radius r=8r=8px, kernel $15\times 15$) obscures alphanumeric content while preserving vehicle context.
@@ C.3 Phase 2: Multi-Agent Orchestration
Three specialized agents coordinate through AutoGen group chat with strict round-robin speaker selection. The Orchestrator Agent manages workflow state and enforces bounded iteration ($n_{\max}=3$). The Auditor Agent performs PII classification via Qwen2.5-VL-32B and conducts independent quality validation of anonymized outputs through visual residual detection. The Generative Agent executes scout-and-zoom segmentation, a two-stage approach motivated by the region proposal paradigm introduced in Faster R-CNN [33], implements IoUIoU-based deduplication ($30\%$ threshold) to prevent redundant reprocessing of overlapping instances, and performs modality-specific inpainting using SDXL with Canny ControlNet for object removal. Complete agent system prompts, tool specifications, PII classification prompts, and inpainting templates are provided in Supplementary. Table 1 details agent responsibilities, tool access, and execution protocols.
Table 1: Agent roles, tool access, and execution protocol in Phase 2 multi-agent workflow. GenerativeAgent implements IoUIoU-based deduplication ($30\%$ threshold) to prevent redundant reprocessing, while AuditorAgent provides independent quality validation.
@@
Protocol: (1) PII Classification: Qwen2.5-VL-32B with structured prompting, returns bboxes in $x_{\min},y_{\min},w,h]$ format. System prompt prohibits acknowledgment-without-execution: “When instructed to use a tool, YOU MUST CALL IT – NEVER just acknowledge.” Performs scout-and-zoom verification for each classified instance: crops to LVLM bbox, runs Grounded-SAM-2 [22, 32], maps mask to full coordinates. (2) Independent Quality Validation: Creates masked image (processed areas blacked out with 5-iteration morphological dilation, $5\times 5$ elliptical kernel), performs LVLM-based residual detection to verify no visible PII remains. This check is independent of GenerativeAgent’s IoUIoU deduplication, providing dual-layer verification.
@@
Protocol: (1) IoUIoU-Based Deduplication: Before processing, computes $IoU(bbox_{i},bbox_{j})$ for each candidate instance against all processed instances. Instances with $IoU\geq 0.3$ skipped with iou_overlap_with_processed status, preventing redundant work when AuditorAgent detects the same region repeatedly (e.g., IoU=$0.88$ in berlin_000002 iteration 2). Threshold config. via PII_IOU_THRESHOLD (default $0.3$). (2) Scout-and-Zoom Segmentation: Crops to LVLM bbox, runs Grounded-SAM-2 on crop, maps precise mask to full coordinates. Failed segmentations skipped with scout_zoom_failed_no_fallback status. (3) Modal-Specific Inpainting: SDXL with Canny ControlNet ($\alpha=0.3$, strength=$0.9$, guidance=$9.0$, $25$ denoising steps) for objects/text using LVLM description as prompt. Person inpainting (SDXL with OpenPose ControlNet, Phase 1) handled deterministically. All color matching disabled (luminance=$0.0$, chrominance=$0.0$) to prevent appearance correlation. (4) Format Enforcement: Handles three LLM serialization errors via robust parsing: single strings with comma-separated dicts ("{a:1},{b:2}"), stringified dict objects (["{a:1}"]), nested list structures. Format examples enforce JSON dict objects.
@@ C.4 Phase 2: Bounded Iterative Refinement
The framework implements bounded PDCA cycles (Figure 2) ensuring coverage while guaranteeing termination. Each cycle executes: (1) Plan–Orchestrator determines which instances require processing; (2) Do–Generative performs IoUIoU deduplication, then executes scout-and-zoom segmentation and inpainting; (3) Check–Generative validates bbox overlap with processed instances (efficiency), then Auditor independently validates visual completeness on masked image (quality); (4) Act–Orchestrator decides whether to continue ($n<n_{\max}$ AND residuals exist) or terminate. The dual-layer validation mechanism balances efficiency and quality: GenerativeAgent’s IoUIoU deduplication prevents redundant computation when AuditorAgent detects the same instance across iterations (skip($bbox_{i}$) = true if $\max_{j}IoU(bbox_{i},bbox_{j})\geq\tau$ where $\tau=0.3$, false otherwise), while AuditorAgent’s independent visual inspection ensures anonymization quality.
@@ D.1.2 Re-ID Risk Assessment
Evaluation Protocol. We adopt a realistic threat model simulating adversarial Re-ID attempts: (1) Train a ResNet50 [10] Re-ID model with triplet loss [34] and center loss [39] on the original CUHK03-NP detected training set ($767$ identities, $7,365$ images) for $120$ epochs; (2) Test using original query images ($1,400$ queries) against anonymized gallery images ($5,332$ gallery) with re-ranking [46]. This simulates an attacker with access to original surveillance footage attempting to re-identify individuals by matching against anonymized public releases. High matching accuracy indicates privacy risk; low accuracy confirms effective anonymization. This threat model targets automated Re-ID at scale, the primary vector for mass surveillance. Acquaintances may still recognize subjects through retained structural features (pose, build), which utility-preserving methods must retain; person Re-ID threat models primarily address automated systems, not determined observers with prior knowledge. We employ standard Re-ID metrics: Rank-1 ($R1$) accuracy measures the percentage of queries where the correct identity appears as the top-ranked match in the gallery, representing the success rate of immediate Re-ID. Mean Average Precision (mAPmAP) evaluates ranking quality across all gallery positions, capturing the overall retrieval performance.
Results. Table 4 demonstrates our method implements substantial Re-ID risk reduction: only $16.9\%$ $R1$ accuracy and $13.7\%$ mAPmAP, representing $72.9\%$ and $79.2\%$ reductions from original performance ($62.4\%$ $R1$, $66.0\%$ mAPmAP), respectively. In comparison to FADM [47], we reduce person Re-ID risk ($R1$: $16.9\%$ vs. $33.4\%$) and (mAPmAP: $13.7\%$ vs. $32.9\%$) while improving distribution preservation (Table 2: $56\%$ better KID, $39\%$ better FID). Aggressive obfuscation methods (DP2: $8.6\%$ $R1$, $4.4\%$ mAPmAP; Gauss. Blur: $9.4\%$ $R1$, $6.4\%$ mAPmAP) achieve lower Re-ID rates but at severe quality cost. DP2 exhibits worse perceptual distortion (LPIPS: $0.303$ vs. $0.203$) and worse distribution distortion (KID: $0.066$ vs. $0.014$). Gauss. Blur shows catastrophic degradation with worse distribution alignment (KID: $0.224$).
@@ D.1.3 PII Detection Quality
For the supervised baseline, we report APAP at IoUIoU thresholds 0.5:0.95. For unified mask comparison: Dice (F1) = $\frac{2|P\cap G|}{|P|+|G|}$, IoU = $\frac{|P\cap G|}{|P\cup G|}$, and pixel-level Prec./Rec., where $P$ is predicted and $G$ is ground truth.
@@ 4. Select top 5 most sensitive/prominent instances
Given an image II and existing mask set $\mathcal{M}$, identify context-dependent PII instances $\mathcal{P}=\{p_{1},\ldots,p_{k}\}$ where each $p_{i}=(\text{desc}_{i},\text{bbox}_{i})$ represents a physical description and spatial location.
@@
anonymize_and_inpaint_persons_tool executes three-stage person anonymization with attribute diversification. First, Qwen2.5-VL-32B extracts semantic attributes $\mathcal{A}_{i}$ (pose, clothing, action) for each person ii. Second, Qwen2.5-32B-Instruct transforms $\mathcal{A}_{i}\rightarrow\mathcal{A}^{\prime}_{i}$ by sampling colors from the diverse palette $\mathcal{C}$, decoupling description from diversification. Third, SDXL with OpenPose ControlNet generates anonymized person $I^{\prime}_{i}$ using pose-preserved synthesis with the configuration above. Color matching is disabled ($\beta=0.0$) to allow complete appearance transformation while OpenPose conditioning maintains pose realism and context consistency. The function returns anonymized count, output path, and processing time.
@@
We train Mask R-CNN [9] with ResNet-50-FPN [20] on Visual Redactions Dataset [27, 28]: 3,8733{,}873 training images, 21,48921{,}489 annotations across $28$ privacy categories. COCO-pretrained weights initialize the backbone. Training uses SGD (momentum $0.9$, weight decay $0.0001$), base LR $0.002$ with warmup ($500$ iterations), decay at 1818K/2525K iterations, batch size 88 ($2\times$ L40S GPUs), 3030K iterations ($\sim$$5.7$ hours). Augmentation: random horizontal flip, multi-scale training (640640800800px shortest edge).

### 03 Apr 2026 v2

@@ Abstract
Street-level imagery contains personally identifiable information (PII), some of which is context-dependent. Existing anonymization methods either over-process images or miss subtle identifiers, while API-based solutions compromise data sovereignty. We present an agentic framework CAIAMAR (Context-Aware Image Anonymization with Multi-Agent Reasoning) for context-aware PII segmentation with diffusion-based anonymization, combining pre-defined processing for high-confidence cases with multi-agent reasoning for indirect identifiers. Three specialized agents coordinate via round-robin speaker selection in a Plan-Do-Check-Act (PDCA) cycle, enabling large vision-language models to classify PII based on spatial context (private vs. public property) rather than rigid category rules. The agents implement spatially-filtered coarse-to-fine detection where a scout-and-zoom strategy identifies candidates, open-vocabulary segmentation processes localized crops, and $IoU$-based deduplication ($30\%$ threshold) prevents redundant processing. Modal-specific diffusion guidance with appearance decorrelation substantially reduces re-identification (Re-ID) risks. On CUHK03-NP, our method reduces person Re-ID risk by $73\%$ ($R1$: $16.9\%$ vs. $62.4\%$ baseline). For image quality preservation on CityScapes, we achieve KID: $0.001$, and FID: $9.1$, significantly outperforming existing anonymization. The agentic workflow detects non-direct PII instances across object categories, and downstream semantic segmentation is preserved. Operating entirely on-premise with open-source models, the framework generates human-interpretable audit trails supporting EU’s GDPR transparency requirements while flagging failed cases for human review.
@@ B.1 Obfuscation and Generative Methods
Conventional image anonymization approaches, such as EgoBlur [30], rely on blurring for efficiency. However, these methods insufficiently protect identity information while degrading downstream task performance. For instance, a $5.3\%$ drop in instance segmentation average precision ($AP$) on CityScapes [5] has been reported when applying Gaussian blur [13]. Moreover, blurring remains vulnerable to inversion attacks: $95.9\%$ identity retrieval accuracy has been achieved on blurred CelebA-HQ [15] face images at $256\times 256$ resolution [43], demonstrating that simple obfuscation is inadequate for modern threat models. Generative methods address these limitations by synthesizing realistic replacements for sensitive regions, preserving contextual consistency while reducing Re-ID risks. Early deep learning–based methods predominantly use Generative Adversarial Networks (GANs). DeepPrivacy [14] employs a conditional GAN that generates anonymized faces conditioned on pose keypoints and background context. DeepPrivacy2 [12] extends this approach to full-body anonymization via a style-based GAN, improving realism and consistency across human appearance and pose. Recent advances leverage diffusion models, offering higher fidelity, controllability, and semantic alignment. LDFA [16] applies latent diffusion for face anonymization in street recordings, maintaining both realism and privacy. The FADM framework [47] and RAD [25] employ diffusion to anonymize persons. SVIA [21] extends diffusion-based anonymization to large-scale street-view imagery, ensuring privacy preservation in autonomous driving datasets by replacing the semantic categories of person, vehicle, traffic sign, road, and building. Reverse Personalization [17] introduces a finetuning-free face anonymization framework based on conditional diffusion inversion with null-identity embeddings. By applying negative classifier-free guidance, it successfully steers the generation process away from identity-defining features while enabling attribute-controllable anonymization for arbitrary subjects.
@@ C.2 Phase 1: Pre-defined Preprocessing
We employ YOLOv8m-seg [^1] https://github.com/ultralytics/ultralytics, Accessed October 25, 2025 for instance segmentation (confidence threshold $\tau=0.25$). The anonymization pipeline operates in three stages: (1) extracts visual descriptions via Qwen2.5-VL-32B [2], capturing pose, body build, view angle, and actions; (2) generates anonymized prompts by instructing the LVLM to randomly select clothing colors and brightness modifiers from predefined palettes ($20$ colors, $10$ brightness levels) [^2] Color palette: gray, beige, navy, white, black, brown, khaki, blue, red, green, yellow, pink, teal, burgundy, olive, charcoal, maroon, tan, cream, sage. Brightness modifiers: light, dark, bright, faded, vibrant, muted, pale, deep, pastel, bold. rather than describing observed clothing, while preserving utility-relevant attributes (body build, pose, view angle, actions) and uniformed personnel context; (3) performs SDXL inpainting at $768$ px resolution with OpenPose [3] ControlNet333thibaud/controlnet-openpose-sdxl-1.0 (conditioning scale $\alpha=0.8$, strength $\rho=0.9$, guidance scale $s=9.0$, $25$ denoising steps). The positive and negative prompts (excluding the LVLM-generated description) are adapted from [47]. Critically, we disable all color matching to prevent appearance correlation and eliminate Re-ID vectors.
License Plate Processing. Small object detection challenges necessitate specialized handling. We train YOLOv8s on UC3M-LP dataset [31] at 12801280px resolution, achieving $mAP_{50-95}=0.82$. At inference, we apply reduced confidence threshold ($\tau=0.05$) prioritizing recall and Non-Maximum Suppression (NMS) with $IoU$ threshold of $0.5$ to ensure individual plate separation, aligning with conventional NMS practices [37]. Gauss. blur (radius r=8r=8px, kernel $15\times 15$) obscures alphanumeric content while preserving vehicle context.
@@ C.3 Phase 2: Multi-Agent Orchestration
Three specialized agents coordinate through AutoGen group chat with strict round-robin speaker selection. The Orchestrator Agent manages workflow state and enforces bounded iteration ($n_{\max}=3$). The Auditor Agent performs PII classification via Qwen2.5-VL-32B and conducts independent quality validation of anonymized outputs through visual residual detection. The Generative Agent executes scout-and-zoom segmentation, a two-stage approach motivated by the region proposal paradigm introduced in Faster R-CNN [33], implements $IoU$-based deduplication ($30\%$ threshold) to prevent redundant reprocessing of overlapping instances, and performs modality-specific inpainting using SDXL with Canny ControlNet for object removal. Complete agent system prompts, tool specifications, PII classification prompts, and inpainting templates are provided in Supplementary. Table 1 details agent responsibilities, tool access, and execution protocols.
Table 1: Agent roles, tool access, and execution protocol in Phase 2 multi-agent workflow. GenerativeAgent implements $IoU$-based deduplication ($30\%$ threshold) to prevent redundant reprocessing, while AuditorAgent provides independent quality validation.
@@
Protocol: (1) PII Classification: Qwen2.5-VL-32B with structured prompting, returns bboxes in $x_{\min},y_{\min},w,h]$ format. System prompt prohibits acknowledgment-without-execution: “When instructed to use a tool, YOU MUST CALL IT – NEVER just acknowledge.” Performs scout-and-zoom verification for each classified instance: crops to LVLM bbox, runs Grounded-SAM-2 [22, 32], maps mask to full coordinates. (2) Independent Quality Validation: Creates masked image (processed areas blacked out with 5-iteration morphological dilation, $5\times 5$ elliptical kernel), performs LVLM-based residual detection to verify no visible PII remains. This check is independent of GenerativeAgent’s $IoU$ deduplication, providing dual-layer verification.
@@
Protocol: (1) $IoU$-Based Deduplication: Before processing, computes $IoU(bbox_{i},bbox_{j})$ for each candidate instance against all processed instances. Instances with $IoU\geq 0.3$ skipped with iou_overlap_with_processed status, preventing redundant work when AuditorAgent detects the same region repeatedly (e.g., IoU=$0.88$ in berlin_000002 iteration 2). Threshold config. via PII_IOU_THRESHOLD (default $0.3$). (2) Scout-and-Zoom Segmentation: Crops to LVLM bbox, runs Grounded-SAM-2 on crop, maps precise mask to full coordinates. Failed segmentations skipped with scout_zoom_failed_no_fallback status. (3) Modal-Specific Inpainting: SDXL with Canny ControlNet ($\alpha=0.3$, strength=$0.9$, guidance=$9.0$, $25$ denoising steps) for objects/text using LVLM description as prompt. Person inpainting (SDXL with OpenPose ControlNet, Phase 1) handled deterministically. All color matching disabled (luminance=$0.0$, chrominance=$0.0$) to prevent appearance correlation. (4) Format Enforcement: Handles three LLM serialization errors via robust parsing: single strings with comma-separated dicts ("{a:1},{b:2}"), stringified dict objects (["{a:1}"]), nested list structures. Format examples enforce JSON dict objects.
@@ C.4 Phase 2: Bounded Iterative Refinement
The framework implements bounded PDCA cycles (Figure 2) ensuring coverage while guaranteeing termination. Each cycle executes: (1) Plan–Orchestrator determines which instances require processing; (2) Do–Generative performs $IoU$ deduplication, then executes scout-and-zoom segmentation and inpainting; (3) Check–Generative validates bbox overlap with processed instances (efficiency), then Auditor independently validates visual completeness on masked image (quality); (4) Act–Orchestrator decides whether to continue ($n<n_{\max}$ AND residuals exist) or terminate. The dual-layer validation mechanism balances efficiency and quality: GenerativeAgent’s $IoU$ deduplication prevents redundant computation when AuditorAgent detects the same instance across iterations (skip($bbox_{i}$) = true if $\max_{j}IoU(bbox_{i},bbox_{j})\geq\tau$ where $\tau=0.3$, false otherwise), while AuditorAgent’s independent visual inspection ensures anonymization quality.
@@ D.1.2 Re-ID Risk Assessment
Evaluation Protocol. We adopt a realistic threat model simulating adversarial Re-ID attempts: (1) Train a ResNet50 [10] Re-ID model with triplet loss [34] and center loss [39] on the original CUHK03-NP detected training set ($767$ identities, $7,365$ images) for $120$ epochs; (2) Test using original query images ($1,400$ queries) against anonymized gallery images ($5,332$ gallery) with re-ranking [46]. This simulates an attacker with access to original surveillance footage attempting to re-identify individuals by matching against anonymized public releases. High matching accuracy indicates privacy risk; low accuracy confirms effective anonymization. This threat model targets automated Re-ID at scale, the primary vector for mass surveillance. Acquaintances may still recognize subjects through retained structural features (pose, build), which utility-preserving methods must retain; person Re-ID threat models primarily address automated systems, not determined observers with prior knowledge. We employ standard Re-ID metrics: Rank-1 ($R1$) accuracy measures the percentage of queries where the correct identity appears as the top-ranked match in the gallery, representing the success rate of immediate Re-ID. Mean Average Precision ($mAP$) evaluates ranking quality across all gallery positions, capturing the overall retrieval performance.
Results. Table 4 demonstrates our method implements substantial Re-ID risk reduction: only $16.9\%$ $R1$ accuracy and $13.7\%$ mAPmAP, representing $72.9\%$ and $79.2\%$ reductions from original performance ($62.4\%$ $R1$, $66.0\%$ mAPmAP), respectively. In comparison to FADM [47], we reduce person Re-ID risk ($R1$: $16.9\%$ vs. $33.4\%$) and ($mAP$: $13.7\%$ vs. $32.9\%$) while improving distribution preservation (Table 2: $56\%$ better KID, $39\%$ better FID). Aggressive obfuscation methods (DP2: $8.6\%$ $R1$, $4.4\%$ mAPmAP; Gauss. Blur: $9.4\%$ $R1$, $6.4\%$ mAPmAP) achieve lower Re-ID rates but at severe quality cost. DP2 exhibits worse perceptual distortion (LPIPS: $0.303$ vs. $0.203$) and worse distribution distortion (KID: $0.066$ vs. $0.014$). Gauss. Blur shows catastrophic degradation with worse distribution alignment (KID: $0.224$).
@@ D.1.3 PII Detection Quality
For the supervised baseline, we report $AP$ at $IoU$ thresholds 0.5:0.95. For unified mask comparison: Dice (F1) = $\frac{2|P\cap G|}{|P|+|G|}$, IoU = $\frac{|P\cap G|}{|P\cup G|}$, and pixel-level Prec./Rec., where $P$ is predicted and $G$ is ground truth.
@@ 4. Select top 5 most sensitive/prominent instances
Given an image $I$ and existing mask set $\mathcal{M}$, identify context-dependent PII instances $\mathcal{P}=\{p_{1},\ldots,p_{k}\}$ where each $p_{i}=(\text{desc}_{i},\text{bbox}_{i})$ represents a physical description and spatial location.
@@
anonymize_and_inpaint_persons_tool executes three-stage person anonymization with attribute diversification. First, Qwen2.5-VL-32B extracts semantic attributes $\mathcal{A}_{i}$ (pose, clothing, action) for each person $i$. Second, Qwen2.5-32B-Instruct transforms $\mathcal{A}_{i}\rightarrow\mathcal{A}^{\prime}_{i}$ by sampling colors from the diverse palette $\mathcal{C}$, decoupling description from diversification. Third, SDXL with OpenPose ControlNet generates anonymized person $I^{\prime}_{i}$ using pose-preserved synthesis with the configuration above. Color matching is disabled ($\beta=0.0$) to allow complete appearance transformation while OpenPose conditioning maintains pose realism and context consistency. The function returns anonymized count, output path, and processing time.
@@
We train Mask R-CNN [9] with ResNet-50-FPN [20] on Visual Redactions Dataset [27, 28]: 3,8733{,}873 training images, 21,48921{,}489 annotations across $28$ privacy categories. COCO-pretrained weights initialize the backbone. Training uses SGD (momentum $0.9$, weight decay $0.0001$), base LR $0.002$ with warmup ($500$ iterations), decay at 1818K/2525K iterations, batch size $8$ ($2\times$ L40S GPUs), 3030K iterations ($\sim$$5.7$ hours). Augmentation: random horizontal flip, multi-scale training (640640800800px shortest edge).